OpenStudy (anonymous):

Which answer shows this equation in standard form? 8 – 4(x – y) = –6x + 6

OpenStudy (godlovesme):

first distribute -4 do u have choices?

OpenStudy (anonymous):

yes 1 sec

OpenStudy (godlovesme):

k :)

OpenStudy (godlovesme):

u there? @No_Flex_Zone

OpenStudy (anonymous):

A. 2x + 4y = 2 B. 2x + 4y = –2 C. 2x – 4y = –2 D. –10x + 4y = –2

OpenStudy (godlovesme):

let's solve this :)

OpenStudy (anonymous):

ok

OpenStudy (godlovesme):

 can u try to get the x's to the left ?

OpenStudy (godlovesme):

 is that one of the options? @No_Flex_Zone

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (godlovesme):

welcome thanks for the medal :)

OpenStudy (anonymous):

yourwelcome

OpenStudy (anonymous):

can u help me with another one

OpenStudy (godlovesme):

sure open a new question and tag me:)

OpenStudy (anonymous):

ok